# Dirac branch-cut modes

### 30 Mar 2026 v1

@@ Abstract
The emergence of bound states in Dirac media is of profound importance for many fields of physics, including particle physics, condensed matter physics, and photonics. For example, the Jackiw–Rebbi state [1], which appears at a domain wall in one-dimensional (1D) space where a Dirac mass field changes sign, is recognized as a theoretical prototype for the edge states of topological insulators in 1D or higher dimensions [2]. Such edge states have many interesting properties, and have been observed in a variety of real materials [3, 4], as well as metamaterials based on photonics, acoustics, and other classical-wave platforms [5, 6, 7, 8, 9, 10]. The concept has also been extended beyond domain walls to features such as higher-order boundaries [11, 12, 13, 14] and spatial defects [15, 16, 17, 18, 19, 20]. Notably, a leading conceptual framework for understanding defect-bound topological states is the Jackiw–Rossi model [21], a generalization of the Jackiw–Rebbi model that features a Dirac field in two-dimensional (2D) space with a complex mass field (composed of two independent Dirac mass parameters); a vortex in the phase of the complex mass binds a localized zero-energy state that is robust against mass field variations that do not alter the winding number [22]. These Jackiw–Rossi states form the conceptual foundation for Majorana excitations in topological superconductors [23, 24], and have recently also been experimentally realized in Kekulé-modulated lattices [25, 26, 27]. In photonics, particularly, there is much ongoing research into using “Dirac vortex” modes for enhancing light-matter interactions, designing new laser cavity designs, and other practical applications [28, 29, 30, 31, 32, 33].
@@
Figure 1: Properties of Dirac branch-cut (DBC) modes. a, Upper panel: complex plane for the Jackiw–Rossi mass parameter $m$, indicating two points with the same $|m|$ and phase difference $\Delta\theta$. Lower panel: a schematic of a piecewise-constant $m(\mathbf{r})$ distribution in 2D space, constructed from those two $m$ values with phase dislocation along $y=0$. b, Phase dislocation mode (DBC mode) energies versus wavenumber $k_{x}$ for $\Delta\theta=\pi$ (left panel) and $\Delta\theta=2\pi/3$ (middle panel), and versus $\Delta\theta$ at $k_{x}=0$ (right panel). Cyan dashes show the results obtained analytically using Eqs. (4)–(5). Color maps show the minimum eigenvalue magnitude of the scattering matrix, computed directly from the Hamiltonian (1); a zero value corresponds to a boundary state (see Supplementary Information). The bulk bands are shaded in gray. c, Phase profile of a mass distribution made from a branch function $m=(z-3)^{1/2}(z+3)^{1/2}$, where $z=x+iy$. d, Numerically-obtained DBC mode energies versus mode order for the mass distribution in c. e, Spatial profiles for two of the modes in d: the Jackiw–Rossi state at $E=0$ (left panel), and the order-4 mode (right panel). f, Phase profile of a “phase-only” mass distribution $m=m_{0}\mathrm{arg}[(z-3)^{1/2}(z+3)^{-1/2}]$, with $m_{0}=5$. g, DBC mode energies for the mass distribution in f (blue markers), and for another mass distribution with $\Delta\theta=2\pi/3$ using the same cut (yellow markers). The theoretically-predicted minigap for the latter is indicated by horizontal dashes. h Representative spatial profiles of two of the DBC modes for $\Delta\theta=\pi$. All the DBC modes have very similar confinement lengths, up to high orders approaching the bulk bands.
@@
In Eq. (4), the effective mass $m_{b}$ depends on $m_{0}$ and the phase dislocation $\Delta\theta$. For $\Delta\theta=\pi$, which corresponds to a Jackiw–Rebbi-like sign flip of $m$ across the domain wall, we have $m_{b}=0$; for other $\Delta\theta$, we have $0<m_{b}<m_{0}$ (i.e., there is a minigap enclosed within the bulk gap). The addition of a phase gradient $v_{\theta}y$ retains this result, but with a correction to $m_{b}$ (see Methods). Importantly, because $\kappa$ is independent of $k_{x}$, all the DBC modes have the same transverse confinement length, regardless of where they lie in the bulk gap. This property is due to the 1D dispersion relation (4) and the bulk dispersion relation for guided modes, $E^{2}=k_{x}^{2}-\kappa^{2}+m_{0}^{2}$, from Eq. (1); these involve the same Dirac velocity, and combining them leads to Eq. (5).
@@
In Fig. 1c, we show the distribution of $\theta(\mathbf{r})=\mathrm{arg}[m(\mathbf{r})]$ for $m(\mathbf{r})=(z-3)^{1/2}(z+3)^{1/2}$. The cut is chosen to be a sinusoidal curve connecting the branch points at $z=\pm 3$, and the phase discontinuity is $\Delta\theta=\pi$. Since $|m|\rightarrow\infty$ as $|\mathbf{r}|\rightarrow\infty$, there are no free solutions, only bound states. We solve for these numerically using the finite element method (FEM) (see Methods), obtaining the spectrum in Fig. 1d, with discrete energies extending to $E\rightarrow\pm\infty$. As the branch function has a phase winding of $2\pi$ upon encircling both branch points, the index theorem for the Jackiw–Rossi model predicts a mode at $E=0$ [21, 22], consistent with our obtained spectrum; in other words, the Jackiw–Rossi state is one of the DBC modes. However, all the DBC modes are strongly bound to the cut, with similar confinement lengths; two of them, the zero mode and the order-4 mode, are plotted in Fig. 1e.
@@ Methods
Starting from the Hamiltonian (1), we let the mass distribution be invariant along $x$, i.e., $m_{1}+im_{2}=m(y)$. We aim to derive the boundary states for a phase dislocation along the line $y=0$. Introducing the wavenumber $k_{x}$, the Schrödinger equation reduces to
@@
When this is plugged into (7), the $\partial_{y}$ gives rise to an additional gauge term because $\theta$ depends implicitly on $y$:
@@
where $\kappa\in\mathbb{R}^{+}$ and $|\chi_{\pm}\rangle$ are four-component spinors. Plugging this into (10), for $y\neq 0$, gives
@@
To handle the boundary conditions at $y=0$, we require the original wavefunction $\psi(y)$ to be continuous at $y=0$, so the rotated wavefunction $\varphi(y)$ is discontinuous. Integrating (10) across an infinitesimal interval over $y=0$ using Eq. (12), we obtain
$-i\tau_{3}\sigma_{2}\left[\varphi(0^{+})-\varphi(0^{-})\right]$ $=\frac{\Delta\theta}{4}\sigma_{2}\left[\varphi(0^{+})+\varphi(0^{-})\right].$
Referring to (13), this is satisfied by the boundary condition $|\chi_{+}\rangle=U_{\Delta\theta}\;|\chi_{-}\rangle$. We can plug this into Eq. (14) to eliminate (say) $\chi_{+}$, obtaining
@@
Using (15), the terms in parentheses can be simplified to
$2i\kappa\tau_{3}\sigma_{2}+m_{0}\left[\tau_{1}(\cos\Delta\theta-1)+\tau_{2}\sin\Delta\theta\right].$
What is the spectrum of this operator? Pick an eigenstate of $\sigma_{2}$, i.e., $|\chi_{-}\rangle=|u\rangle\otimes|v\rangle$ where $\sigma_{2}|v\rangle=\pm|v\rangle$. Then (16) reduces to
@@
Note that the result for $\kappa$ solves (17) for both choices of $\pm$, i.e., both eigenstates of $\sigma_{2}$. It follows that $|\chi_{-}\rangle$ is a superposition of these two choices. To be specific, let us take the eigenstates of $\sigma_{2}$ to be $|v_{+}\rangle=2^{-1/2}[1,i]^{T}$ and $|v_{-}\rangle=\sigma_{1}|v_{+}\rangle$. In each case, the non-Hermitian matrix in (17) has equal and opposite eigenvalues that are either purely real or imaginary. The zero eigenvalue condition is an exceptional point where the matrix has only one eigenvector, which we denote by $|u_{+}\rangle$ or $|u_{-}\rangle$. Evidently, $|u_{+}\rangle$ and $|u_{-}\rangle$ depend only on $\Delta\theta$ (and are independent of $m_{0}$ and $k_{x}$), and can be defined such that
@@
where $\kappa$ is given by (5) and
@@
This 1D Dirac equation, with effective Dirac mass $m_{b}$, gives the dispersion relation (4). The spectral theorem now implies that the restricted eigenproblem (20) has two solutions for each $k_{x}$, which is significant as restricted eigenproblems are not guaranteed to possess solutions. This clarifies why the $4\times 4$ Dirac equation exhibits two, not four, boundary states for each $k_{x}$, and retroactively justifies the ansatz (13).
@@ Acknowledgements
Supplemental Materials for
@@
The acoustic structure that we use to realize the Jackiw–Rossi model, and the consequent DBC modes, is similar to the design in Ref. 44. We begin with a triangular lattice of solid pillars surrounded by air, as shown in Fig. 2a of the main text. If all pillars have equal radii, the band structure has two Dirac points at the corners of the hexagonal Brillouin zone. If the lattice is viewed in terms of an expanded unit cell with six pillars each, both Dirac points are folded to the center, $\Gamma$, of the shrunken Brillouin zone [8]. In Fig. S3a, the gray lines show the acoustic band structure calculated by FEM. Near $\Gamma$, therefore, we have four bands (two from each valley) governed by an effective Jackiw–Rossi model with $m_{1}=m_{2}=0$.

### 02 Apr 2026 v2

@@ Abstract
The emergence of bound states in Dirac media is of profound importance for many fields of physics, including particle physics, condensed matter physics, and photonics. For example, the Jackiw–Rebbi state [1], which appears at a domain wall in one-dimensional (1D) space where a Dirac mass field changes sign, is recognized as a theoretical prototype for the edge states of topological insulators in 1D or higher dimensions [2]. Such edge states have many interesting properties, and have been observed in a variety of real materials [3, 4], as well as metamaterials based on photonics, acoustics, and other classical-wave platforms [5, 6, 7, 8, 9, 10]. The concept has also been extended beyond domain walls to features such as higher-order boundaries [11, 12, 13, 14, 15, 16, 17, 18] and spatial defects [19, 20, 21, 22, 23, 24]. Notably, a leading conceptual framework for understanding defect-bound topological states is the Jackiw–Rossi model [25], a generalization of the Jackiw–Rebbi model that features a Dirac field in two-dimensional (2D) space with a complex mass field (composed of two independent Dirac mass parameters); a vortex in the phase of the complex mass binds a localized zero-energy state that is robust against mass field variations that do not alter the winding number [26]. These Jackiw–Rossi states form the conceptual foundation for Majorana excitations in topological superconductors [27, 28, 29], and have recently also been experimentally realized in Kekulé-modulated lattices [30, 31, 32]. In photonics, particularly, there is much ongoing research into using “Dirac vortex” modes for enhancing light-matter interactions, designing new laser cavity designs, and other practical applications [33, 34, 35, 36, 37, 38].
@@
Figure 1: Properties of Dirac branch-cut (DBC) modes. a, Upper panel: complex plane for the Jackiw–Rossi mass parameter $m$, indicating two points with the same $|m|$ and phase difference $\Delta\theta$. Lower panel: a schematic of a piecewise-constant $m(\mathbf{r})$ distribution in 2D space, constructed from those two $m$ values with phase dislocation along $y=0$. b, Phase dislocation mode (DBC mode) energies versus wavenumber $k_{x}$ for $\Delta\theta=\pi$ (left panel) and $\Delta\theta=2\pi/3$ (middle panel), and versus $\Delta\theta$ at $k_{x}=0$ (right panel). Cyan dashes show the results obtained analytically using Eqs. (4)–(5). Color maps show the minimum eigenvalue magnitude of the scattering matrix, computed directly from the Hamiltonian (1); a zero value corresponds to a boundary state (see Supplementary Information). The bulk bands are shaded in gray. c, Phase profile of a mass distribution made from a branch function $m=(z-3)^{1/2}(z+3)^{1/2}$, where $z=x+iy$. d, Numerically-obtained DBC mode energies versus mode order with the phase of mass distribution in c, and the amplitude extracted from the branch function $|m|$. e, Spatial profiles for two of the modes in d: the Jackiw–Rossi state at $E=0$ (left panel), and the order-4 mode (right panel). f, Phase profile of a “phase-only” mass distribution $m=m_{0}\mathrm{arg}[(z-3)^{1/2}(z+3)^{-1/2}]$, with $m_{0}=5$. g, DBC mode energies for the mass distribution in f (blue markers), and for another mass distribution with $\Delta\theta=2\pi/3$ using the same cut (yellow markers). The theoretically-predicted minigap for the latter is indicated by horizontal dashes. h Representative spatial profiles of two of the DBC modes for $\Delta\theta=\pi$. For the DBC mode profiles for $\Delta\theta=2\pi/3$, see Supplementary Information Figs. S1 and S2. All the DBC modes have very similar confinement lengths, up to high orders approaching the bulk bands.
@@
In Eq. (4), the effective mass $m_{b}$ depends on $m_{0}$ and the phase dislocation $\Delta\theta$. For $\Delta\theta=\pi$, which corresponds to a Jackiw–Rebbi-like sign flip of $m$ across the domain wall, we have $m_{b}=0$; for other $\Delta\theta$, we have $0<m_{b}<m_{0}$ (i.e., there is a minigap enclosed within the bulk gap). The addition of a phase gradient $v_{\theta}y$ retains this result, but with a correction to $m_{b}$ (see Methods). Importantly, because $\kappa$ is independent of $k_{x}$, all the DBC modes have the same transverse confinement length, regardless of where they lie in the bulk gap. This property is due to the 1D dispersion relation (4) and the bulk dispersion relation for guided modes, $E^{2}=k_{x}^{2}-\kappa^{2}+m_{0}^{2}$, from Eq. (1); these involve the same Dirac velocity, and combining them leads to Eq. (5). Note that this persistent confinement is still dependent on $\Delta\theta$, see Supplementary Information Fig. S1.
@@
In Fig. 1c, we show the distribution of $\theta(\mathbf{r})=\mathrm{arg}[m(\mathbf{r})]$ for $m(\mathbf{r})=(z-3)^{1/2}(z+3)^{1/2}$. The cut is chosen to be a sinusoidal curve connecting the branch points at $z=\pm 3$, and the phase discontinuity is $\Delta\theta=\pi$. Since $|m|\rightarrow\infty$ as $|\mathbf{r}|\rightarrow\infty$, there are no free solutions, only bound states. We solve for these numerically using the finite element method (FEM) (see Methods), based on both the phase and amplitude of mass distributions. The obtained spectrum is shown in Fig. 1d, with discrete energies extending to $E\rightarrow\pm\infty$. As the branch function has a phase winding of $2\pi$ upon encircling both branch points, the index theorem for the Jackiw–Rossi model predicts a mode at $E=0$ [25, 26], consistent with our obtained spectrum; in other words, the Jackiw–Rossi state is one of the DBC modes. However, all the DBC modes are strongly bound to the cut, with similar confinement lengths; two of them, the zero mode and the order-4 mode, are plotted in Fig. 1e. The slightly larger decay lengths near the ends of order-4 mode are due to locally minimum amplitude of branch function.
@@ Methods
Starting from the Hamiltonian Eq. (1), we let the mass distribution be invariant along $x$, i.e., $m_{1}+im_{2}=m(y)$. We aim to derive the boundary states for a phase dislocation along the line $y=0$. Introducing the wavenumber $k_{x}$, the Schrödinger equation reduces to
@@
When this is plugged into Eq. (7), the $\partial_{y}$ gives rise to an additional gauge term because $\theta$ depends implicitly on $y$:
@@
where $\kappa\in\mathbb{R}^{+}$ and $|\chi_{\pm}\rangle$ are four-component spinors. Plugging this into Eq. (10), for $y\neq 0$, gives
@@
To handle the boundary conditions at $y=0$, we require the original wavefunction $\psi(y)$ to be continuous at $y=0$, so the rotated wavefunction $\varphi(y)$ is discontinuous. Integrating Eq. (10) across an infinitesimal interval over $y=0$ using Eq. (12), we obtain
$-i\tau_{3}\sigma_{2}\left[\varphi(0^{+})-\varphi(0^{-})\right]$ $=\frac{\Delta\theta}{4}\sigma_{2}\left[\varphi(0^{+})+\varphi(0^{-})\right].$ (16)
Referring to Eq. (13), this is satisfied by the boundary condition $|\chi_{+}\rangle=U_{\Delta\theta}\;|\chi_{-}\rangle$. We can plug this into Eq. (14) to eliminate (say) $\chi_{+}$, obtaining
@@
Using Eq. (15), the terms in parentheses can be simplified to
$2i\kappa\tau_{3}\sigma_{2}+m_{0}\left[\tau_{1}(\cos\Delta\theta-1)+\tau_{2}\sin\Delta\theta\right].$ (18)
What is the spectrum of this operator? Pick an eigenstate of $\sigma_{2}$, i.e., $|\chi_{-}\rangle=|u\rangle\otimes|v\rangle$ where $\sigma_{2}|v\rangle=\pm|v\rangle$. Then Eq. (17) reduces to
@@
Note that the result for $\kappa$ solves Eq. (19) for both choices of $\pm$, i.e., both eigenstates of $\sigma_{2}$. It follows that $|\chi_{-}\rangle$ is a superposition of these two choices. To be specific, let us take the eigenstates of $\sigma_{2}$ to be $|v_{+}\rangle=2^{-1/2}[1,i]^{T}$ and $|v_{-}\rangle=\sigma_{1}|v_{+}\rangle$. In each case, the non-Hermitian matrix in Eq. (19) has equal and opposite eigenvalues that are either purely real or imaginary. The zero eigenvalue condition is an exceptional point where the matrix has only one eigenvector, which we denote by $|u_{+}\rangle$ or $|u_{-}\rangle$. Evidently, $|u_{+}\rangle$ and $|u_{-}\rangle$ depend only on $\Delta\theta$ (and are independent of $m_{0}$ and $k_{x}$), and can be defined such that
@@
where $\kappa$ is given by Eq. (5) and
@@
This 1D Dirac equation, with effective Dirac mass $m_{b}$, gives the dispersion relation Eq. (4). The spectral theorem now implies that the restricted eigenproblem Eq. (22) has two solutions for each $k_{x}$, which is significant as restricted eigenproblems are not guaranteed to possess solutions. This clarifies why the $4\times 4$ Dirac equation exhibits two, not four, boundary states for each $k_{x}$, and retroactively justifies the ansatz Eq. (13).
@@ Acknowledgements
B.Z. and Y.D.C. conceived the idea. B.Z. performed the theoretical derivations with the help from and Y.D.C. B.Z. performed the numerical calculations and designed the experiments. C.M. conducted the measurements with the help from X.H.Z.. B.Z. analysed the experimental data with the help from C.M. B.Z., Y.D.C. and B.L.Z. wrote the manuscript. Q.W., G.G.L. discussed and commented on the manuscript. Q.J.W., B.L.Z. and Y.D.C. supervised the project.
Supplemental Information for
@@
The acoustic structure that we use to realize the Jackiw–Rossi model, and the consequent DBC modes, is similar to the design in Ref. 49. We begin with a triangular lattice of solid pillars surrounded by air, as shown in Fig. 2a of the main text. If all pillars have equal radii, the band structure has two Dirac points at the corners of the hexagonal Brillouin zone. If the lattice is viewed in terms of an expanded unit cell with six pillars each, both Dirac points are folded to the center, $\Gamma$, of the shrunken Brillouin zone [8]. In Fig. S3a, the gray lines show the acoustic band structure calculated by FEM. Near $\Gamma$, therefore, we have four bands (two from each valley) governed by an effective Jackiw–Rossi model with $m_{1}=m_{2}=0$.